Suppose National Collegiate Athletic Association [NCAA] rules state all student-athletes are to receive an average of 50 hours of academic support, per term. A random sample of 49 University of Maryland student-athletes revealed a mean of 47.5 hours of academic support per term. If the calculated value for the associated test statistic equaled -1.75, what was the standard deviation of the number of hours of academic support the student-athletes in the sample received per term?
Answer #1

Solution

Given that,

mean = \mu = 50

\bar x = 47.5

n = 49

z = -1.75

Using z-score formula  

\bar x = z * \sigma \bar x + \mu \bar x  

47.5 = -1.75 *\sigma\bar x + 50

\sigma\bar x  = 47.5 - 50 / -1.75

\sigma\bar x  = 1.43

standard deviation = \sigma = \sigma \bar x * \sqrt{} n

standard deviation = \sigma = 1.43 * \sqrt{} 49

standard deviation = \sigma = 10